OpenStudy (anonymous):

A population of flies grows according to the function p(x) = 6(3)x, where x is measured in weeks. A local spider has set up shop and consumes flies according to the function s(x) = 4x + 1. What is the population of flies after two weeks with the introduced spider? a) 9 flies b) 45 flies c) 54 flies d) 67 flies

OpenStudy (anonymous):

is this \(p(x)=6\times 3^x\) ?

OpenStudy (anonymous):

my guess is you are supposed to compute \[6\times 3^2-(4\times 2+1)\]

OpenStudy (anonymous):

satellite 73 is correct

OpenStudy (anonymous):

So..basically 45?